OpenStudy (anonymous):

Solve the equation or formula for the variable specified: B=3/7(A-10) , for A

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

B = 3/7(A - 10) -- distribute the 3/7 through the parenthesis B = 3/7A - 30/7 -- multiply the equation by 7 7B = 3A - 30 -- add 30 to each side 7B + 30 = 3A -- divide by 3 (7B + 30)/3 = A or (7/3)B + 10 = A